A planet is revolving around the sun in an elliptical orbit and the distance of planet from sun are 7*10pow15, and 9*10POW15 when is at position p, q. If the speed of planet at p is 6.3*10pow15 its speed at q is

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
Use
I = constant
m* r* w² = const
since w = v/r
Hencw
r* v²/r² = constant
V²/r = const
You will definitely get your answer.
 
Regards
Arun(askIITians forum expert)